Limiting Substance: The substance that is totally reacted upon in a given chemical reaction:

         Mg(OH)2 + 2HCl  → MgCl2 + 2H2O


A chemist has 203 g of Mg(OH)2 and 164 g HCl. The chemist wants to make these substances react to form MgCl2 which can be used for fireproofing wood or as disinfectant. How much MgCl2 can be produced and how much water is formed?


1
Expert's answer
2021-01-13T13:53:06-0500

i) - "x = \frac{4.5}{2} = 2.25" mol Mgcl2

Wmgcl2 = "2.25\times 95"

= 213.75 g

ii)- "x = \frac{4.5\times 2}{2} = 4.5" mol

WH2O = "4.5\times18 = 81g"